# Logarithmic delocalization of low temperature 3D Ising and Potts interfaces above a hard floor

### v1

@@ 3.2. Lower bounding the probability of positivity of the blue interface
Finally, note that $\mathcal{A}_{x,h}^{\mathsf{nonR}}$ was defined in [5, Def. 5.1] as the event of a $\widehat{\mathcal{V}}_{\mathsf{red}}^{c}$-path from $x$ to height $h-1/2$ using only vertices of a random set $\mathcal{P}_{x}$ (which was defined as a random cluster pillar), but this random set $\mathcal{P}_{x}$ is always a subset of $\mathcal{L}_{>0}$ that contains $\mathcal{P}_{x}^{\mathsf{nonR}}$, so restricting to $\mathcal{P}_{x}$ is the same as restricting to $\mathcal{L}_{>0}$ when looking at the connected component of $\widehat{\mathcal{V}}_{\mathsf{red}}^{c}$ vertices containing $x$. ∎
By [5, Prop. 2.24] with the observation that the random cluster pillar $\mathcal{P}_{x}$ defined there necessarily contains $\mathcal{P}_{x}^{\mathsf{nonR}}$, we have the following weak bound on the height of $\mathcal{P}_{x}^{\mathsf{nonR}}$.
@@
There exists $C$ such that for all $x,y$ such that $d(x,\partial\Lambda^{\prime}_{n,\infty})\wedge d(y,\partial\Lambda^{\prime}_{m,\infty})\geq r$,
@@
where the equality used the facts that $h\leq O(\log n)$ and $\alpha_{h}\leq 4(\beta+C)h$ per [5, Proposition 1.5] (a simple forcing argument to lower bound the probability of a deviation via the simplest straight column pillar). Hence, using the inequality $1-x\geq e^{-x/(1-x)}$ for $10<x<$ 1, we get
@@ 3.3. Large deviations upper bound on number of sites not at $h_{n}^{*}$
In $\mathcal{I}_{\mathsf{full}}$, there is a 1-1 correspondence between interfaces and admissible families of walls (quotiented out by vertical translations), where two walls are called admissible if their projections are not 0-connected.
@@
The excess area between two interfaces $\mathcal{I}$ and $\mathcal{J}$ is the difference in their number of plaquettes, $\mathfrak{m}(\mathcal{I};\mathcal{J}):=|\mathcal{I}|-|\mathcal{J}|$. The excess area of a wall $W=(A,B)$ of $\mathcal{I}_{\mathsf{full}}$, is $\mathfrak{m}(W):=|A|-\rho(A)$.
@@
This combinatorial construction (along with some further constructions we do not need in what follows) were used in [14] to establish rigidity of $\mathcal{I}_{\mathsf{full}}$ under $\mu_{n}^{\mathsf{dob}}$. In particular, Peierls-type maps based on those in [6] were used to delete walls from the $1$-to-$1$ representation of interfaces, to get exponential tail bounds on the excess areas of walls of $\mathcal{I}_{\mathsf{full}}$. The following is an immediate corollary of [14, Lem. 15] (see also [15, Lem. 7.132]).
@@
The first line used (3.7) for the bounds on the sums, the binomial is for the choice of $N$ points at which to root the walls of $\mathfrak{U}_{k}$, and the factor of $C^{r}$ is for the number of possible ways to allocate excess area $r$ to the different root points and enumerate over the realizations of the walls. Using the bound $\sum_{j\leq pA}\binom{A}{j}\leq\exp(\mathsf{H}(p)A)$ where $\mathsf{H}(p)$ is the binary entropy function $-p\log p-(1-p)\log(1-p)$, we get
@@
To conclude the upper bound of Theorem 1.1, we note that |$|\{x:\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{blue}})\geq h_{n}^{*}+1\}|\geq\tfrac{C}{\sqrt{\beta}}n^{2}$ is a monotone event in the set of $\mathsf{blue}$ vertices (because it is monotone in the augmented $\mathsf{blue}$ vertex set $\widehat{\mathcal{V}}_{\mathsf{blue}}$, and in turn $\mathcal{V}_{\mathsf{blue}}$ which is expressed as a union of vertices connected by $\mathsf{blue}$ paths to $\partial_{\mathsf{blue}}\Lambda^{\prime}_{n,\infty}$), and thus Proposition 2.9 gives the same for $\mu_{n}^{\mathsf{fl}}$.
@@ 4.1. Comparison of pillar rate to point-to-plane connectivity rate
The main goal of this section is to establish the following.
@@
Finally, for $\mathcal{U}$ finite, let $Z^{\eta}(\mathcal{U})$ denote the partition function for the random cluster measure on $\mathcal{U}$ with boundary conditions $\eta$.
@@
and (letting $\mathcal{V}_{\mathsf{red}}$ under $\mu^{\mathsf{red}}_{n}$ be the connected component of $\mathsf{red}$ vertices of the boundary, analagously to its definition under $\mu^{\mathsf{dob}}_{n}$)
$|\mu_{\infty}^{\mathsf{red}}\big{(}\mathsf{o}\longleftrightarrow\mathcal{L}_{h}\text{ in $\mathcal{V}_{\mathsf{red}}^{c}$}\big{)}-\mu_{n}^{\mathsf{red}}\big{(}\mathsf{o}\longleftrightarrow\mathcal{L}_{h}\text{ in $\mathcal{V}_{\mathsf{red}}^{c}$}\big{)}|=o(1)\,,$
@@
The discrepancy of $2\beta-\log(q-1)$ occurs because of differences of the marginal on $\mathsf{o}=(\frac{1}{2},\frac{1}{2},\frac{1}{2})$, so we will isolate this by defining $\mathcal{E}^{\mathsf{dob}}$ and $\mathcal{E}^{\mathsf{w}}$ as purely random cluster events, intending to capture the typical behavior in the random cluster marginals $\bar{\pi}^{\mathsf{dob}}_{n}$ and $\pi^{\mathsf{w}}_{n}$ when $\sigma_{\mathsf{o}}$ is in $\mathcal{V}_{\mathsf{red}}^{c}$ under ${\mathbf{P}}^{\mathsf{dob}}_{n}$ and ${\mathbf{P}}^{\mathsf{red}}_{n}$ respectively. First, we define the following cones:
@@
The second (more complicated) of the bounds in (4.2) is a consequence of the “isolated pillar” results of [5]. There, a “top pillar” $\mathcal{P}^{\mathsf{top}}_{\mathsf{o}}$ was defined analogous to Definition 3.3 but using $\mathcal{V}_{\mathsf{top}}$ instead of $\mathcal{V}_{\mathsf{red}}$ (we will not use this definition except to reference the results proved in that paper). If $E_{h}$ is the event that the top pillar reaches height $h$, then [5, Thm. 3.8] proves that
@@
We now argue that Eq. 4.2, follows from a simplification of the same arguments of [5] that yielded Eqs. 4.3 and 4.4. The proof of Eq. 4.3 began by defining a cut point of as a vertex $v\in\mathcal{P}^{\mathsf{top}}_{\mathsf{o}}$ with no other vertices in $\mathcal{P}^{\mathsf{top}}_{\mathsf{o}}$ having the same height, and such that removing the four vertical plaquettes surrounding $v$ cuts $\mathcal{C}$ into two 1-connected components, with the component above $v$ contained in $\mathcal{L}_{>\operatorname{ht}(v)}$. The portion of $\mathcal{C}$ above the lowest cut point was labeled the “spine” $\mathcal{S}_{\mathsf{o}}$. There was then a map which straightens the spine when it is too large, and deletes walls in the rest of $\mathcal{C}$ when they are too close to the spine. Doing so deletes the portion of $\mathcal{P}^{\mathsf{top}}_{\mathsf{o}}$ below the spine, so the map finishes by appending a column to rejoin the spine to the rest of $\mathcal{C}$. The proof follows the same way, treating all of $\mathcal{C}$ as the top pillar. Namely, one can define a cut point and the spine of the component $\mathcal{C}$, and apply the same map to the spine portion. Since under $\pi^{\mathsf{w}}_{n}$ there is no need to preserve the disconnection event ${\mathfrak{D}}_{n}$ as there was under $\bar{\pi}^{\mathsf{dob}}_{n}$, the remainder $\mathcal{C}\setminus\mathcal{S}_{\mathsf{o}}$ is handled by the simpler Peierls map which deletes it entirely and replaces it by a column of height $\operatorname{ht}(v)$ up from $\mathsf{o}$. Finally, Eq. 4.4 was shown by observing that when a portion of the spine is straightened, it becomes a column of vertices connected by open edges, and hence under the FK–Potts coupling they are all the same color. Clearly this observation holds regardless of whether the joint measure is ${\mathbf{P}}_{n}^{\mathsf{red}}$ or ${\mathbf{P}}_{n}^{\mathsf{dob}}$.
@@
We begin with (4.5). It is easy to see (by, say a Peierls type argument) that $\pi^{\mathsf{w}}_{n}(\mathcal{E}^{\mathsf{w}})$ is dominated by (i.e., $1-\varepsilon_{\beta}$ of its mass is on) the event that $\mathcal{C}$ is just the six plaquettes surrounding $\mathsf{o}$, which happens with probability $qe^{-6\beta}(1\pm\varepsilon_{\beta})$. Moreover, by rigidity of $\mathcal{I}_{\mathsf{full}}$ (Theorem 2.5)
@@
We turn to establishing (4.6). First note that on $\mathcal{E}^{\mathsf{dob}}$, the event $\operatorname{ht}(\mathcal{P}_{\mathsf{o}}^{\mathsf{nonR}})\geq h$ is equivalent to $A_{\mathsf{o},h}$. Now, intending to utilize the FK–Potts coupling, we wish to first reveal in the random cluster model the component $\mathcal{C}$. In order for the event $A_{\mathsf{o},h}$ to be possible, $\mathcal{C}$ must be a set of plaquettes that confines a chain of vertices connecting $\mathsf{o}$ to height $h$. Crucially, the events $\mathcal{E}^{\mathsf{w}}$ and $\mathcal{E}^{\mathsf{dob}}$ were defined so that the set of plaquettes $C^{*}$ such that $\pi^{\mathsf{w}}_{n}(\mathcal{C}^{*}=C^{*}\mid\mathcal{E}^{\mathsf{w}})$ and $\bar{\pi}^{\mathsf{dob}}_{n}(\mathcal{C}^{*}=C^{*}\mid\mathcal{E}^{\mathsf{dob}})$ both have positive probability is the same, let this set be denoted $\mathbf{C}^{*}$. In particular, $C^{*}\in\mathbf{C}^{*}$ if $C^{*}$ is a 1-connected component of plaquettes in $\mathbf{C}_{\vee}$ that forms part of a bounding surface in the sense that there can be no path in $\mathcal{L}_{>0}$ from $\mathsf{o}$ to $\partial\Lambda_{n,n}^{\prime}$ which does not cross a plaquette of $C^{*}\cup C_{1}$. Thus, we can express
@@
For the denominator, the event $\mathcal{E}^{\mathsf{dob}}$ is equivalent to asking that $\mathcal{C}^{*}\in\mathbf{C}^{*}$, and that the set of plaquettes $\mathcal{C}\setminus\mathcal{C}^{*}$ satisfies a set of criterion independent of $\mathcal{C}^{*}$. (Namely, that $\mathcal{C}\setminus\mathcal{C}^{*}$ contains $C_{1}$, is a subset of $\mathbf{C}_{\curlyvee}\cup C_{1}$, does not cut off $\mathsf{o}$ from $\partial_{\mathsf{blue}}\Lambda_{n,n}^{\prime}$, and combines with $\mathcal{C}^{*}$ such that $\mathcal{C}$ is an interface satisfying ${\mathfrak{D}}$. This last event appears as though the choice of $\mathcal{C}\setminus\mathcal{C}^{*}$ affects the possible choices of $\mathcal{C}^{*}$, but the cut-point at $\mathsf{o}$ means it does not.) Exposing $\mathcal{C}\setminus\mathcal{C}^{*}$ reveals a set of dual-to-open plaquettes around it by maximality. Let $F^{\mathsf{dob}}$ be the union of these plaquettes with $\mathcal{C}\setminus\mathcal{C}^{*}$. Noting that the separation event ${\mathfrak{D}}$ is included in $\mathcal{E}^{\mathsf{dob}}$, we can write $\bar{\pi}^{\mathsf{dob}}_{n}(\cdot\mid\mathcal{E}^{\mathsf{dob}})$ as $\mathbb{E}_{\mathcal{C}\setminus\mathcal{C}^{*}}[\pi^{\gamma}_{\Lambda_{n,n}^{\prime}\setminus F^{\mathsf{dob}}}(\cdot\mid\mathcal{C}^{*}\in\mathbf{C}^{*})\mid\mathcal{E}^{\mathsf{dob}}]$ where $\gamma$ is the boundary condition induced by revealing $\mathcal{C}\setminus\mathcal{C}^{*}$.

### v2

@@ 3.2. Lower bounding the probability of positivity of the blue interface
Finally, note that $\mathcal{A}_{x,h}^{\mathsf{nonR}}$ was defined in [5, Def. 5.1] as the event that there exists a $\widehat{\mathcal{V}}_{\mathsf{red}}^{c}$-path from $x$ to height h1/2h-1/2 using only vertices of a random set $\mathcal{P}_{x}$ (which was defined as a random-cluster pillar), but this random set $\mathcal{P}_{x}$ is always a subset of $\mathcal{L}_{>0}$ that contains $\mathcal{P}_{x}^{\mathsf{nonR}}$, so restricting to $\mathcal{P}_{x}$ is the same as restricting to $\mathcal{L}_{>0}$ when looking at the connected component of $\widehat{\mathcal{V}}_{\mathsf{red}}^{c}$ vertices containing $x$. ∎
By [5, Prop. 2.24] with the observation that the random-cluster pillar $\mathcal{P}_{x}$ defined there necessarily contains $\mathcal{P}_{x}^{\mathsf{nonR}}$, we have the following weak bound on the height of $\mathcal{P}_{x}^{\mathsf{nonR}}$.
@@
There exists a constant $C$ such that for all $x,y$ so that $d(x,\partial\Lambda^{\prime}_{n,\infty})\wedge d(y,\partial\Lambda^{\prime}_{m,\infty})\geq r$,
@@
where the equality used the facts that $h\leq O(\log n)$ and $\alpha_{h}\leq 4(\beta+C)h$ per [5, Proposition 1.5] (a simple forcing argument to lower bound the probability of a deviation via the simplest straight column pillar). Hence, using the inequality $1-x\geq e^{-x/(1-x)}$ for $0<x<1$, we get
@@ 3.3. Large deviations upper bound on number of sites not at $h_{n}^{*}$
In $\mathcal{I}_{\mathsf{full}}$, there is a 1-1 correspondence between interfaces and admissible families of walls (quotiented out by vertical translations), where two walls are called admissible if their projections are not 0-connected and they are contained in the domain (see items (i)–(ii) on page 211 of [15] for the formal statement).
@@
The excess area between two interfaces $\mathcal{I}$ and $\mathcal{J}$ is the difference in their number of plaquettes, $\mathfrak{m}(\mathcal{I};\mathcal{J}):=|\mathcal{I}|-|\mathcal{J}|$. The excess area of a wall $W=(A,B)$ of $\mathcal{I}_{\mathsf{full}}$, is $\mathfrak{m}(W):=|A|-|\rho(W)|:=|A|-|\rho(A\cup B)|$. (When we write $|\rho(F)|$ for a set of faces $F$, we mean the area, or number of faces in $\rho(F)$, so projections of vertical faces do not contribute.)
@@
This combinatorial construction (along with some further constructions we do not need in what follows) were used in [14] to establish rigidity of $\mathcal{I}_{\mathsf{full}}$ under $\mu_{n}^{\mathsf{dob}}$. In particular, Peierls-type maps based on those in [6] were used to delete walls from the 11-to-11 representation of interfaces, to get exponential tail bounds on the excess areas of walls of $\mathcal{I}_{\mathsf{full}}$. The following is an immediate corollary of [14, Lem. 15] (see also [15, Lem. 7.132]).
@@
The first line used (3.7) for the bounds on the sums, the binomial is for the choice of $N$ points at which to root the walls of $\mathfrak{U}_{k}$, and the factor of $C^{r}$ is for the number of possible ways to allocate excess area $r$ to the different root points and enumerate over the realizations of the walls. Using the bound $\sum_{j\leq pA}\binom{A}{j}\leq\exp(\mathsf{H}(p)A)$ where $\mathsf{H}(p)$ is the binary entropy function $p\log p-(1-p)\log(1-p)$, we get
@@
To conclude the upper bound of Theorem 1.1, we note that $|\{x:\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{blue}})\geq h_{n}^{*}+1\}|\geq\tfrac{C}{\sqrt{\beta}}n^{2}$ is a monotone event in the set of $\mathsf{blue}$ vertices (because it is monotone in the augmented $\mathsf{blue}$ vertex set $\widehat{\mathcal{V}}_{\mathsf{blue}}$, and in turn $\mathcal{V}_{\mathsf{blue}}$ which is expressed as a union of vertices connected by $\mathsf{blue}$ paths to $\partial_{\mathsf{blue}}\Lambda^{\prime}_{n,\infty}$), and thus Proposition 2.9 gives the same for $\mu_{n}^{\mathsf{fl}}$.
@@ 4.1. Comparison of pillar rate to point-to-plane connectivity rate
For simplicity, define $A_{\mathsf{o},h}=\{\mathsf{o}\longleftrightarrow\mathcal{L}_{h}\text{ in $\mathcal{V}_{\mathsf{red}}^{c}$}\}$. The main goal of this section is to establish the following.
@@
Finally, for $\mathcal{U}$ finite, let $Z^{\eta}(\mathcal{U})$ denote the partition function for the random-cluster measure on $\mathcal{U}$ with boundary conditions $\eta$.
@@
and
$|\mu_{\infty}^{\mathsf{red}}\big{(}A_{\mathsf{o},h}\big{)}-\mu_{n}^{\mathsf{red}}\big{(}A_{\mathsf{o},h}\big{)}|=o(1)\,$,
@@
The discrepancy of $2\beta-\log(q-1)$ occurs because of differences of the marginal on $\mathsf{o}=(\frac{1}{2},\frac{1}{2},\frac{1}{2})$, so we will isolate this by defining $\mathcal{E}^{\mathsf{dob}}$ and $\mathcal{E}^{\mathsf{w}}$ as purely random-cluster events, intending to capture the typical behavior in the random-cluster marginals $\bar{\pi}^{\mathsf{dob}}_{n}$ and $\pi^{\mathsf{w}}_{n}$ when $\sigma_{\mathsf{o}}$ is in $\mathcal{V}_{\mathsf{red}}^{c}$ under ${\mathbf{P}}^{\mathsf{dob}}_{n}$ and ${\mathbf{P}}^{\mathsf{red}}_{n}$ respectively. First, we define the following cones:
@@
The second (more complicated) of the bounds in (4.2) is a consequence of the “isolated pillar” results of [5]. There, a “top pillar” $\mathcal{P}^{\mathsf{top}}_{\mathsf{o}}$ was defined analogous to Definition 3.3 but using $\mathcal{V}_{\mathsf{top}}$ instead of $\mathcal{V}_{\mathsf{red}}$ (we will not use this definition except to reference the results proved in that paper). If $E_{h}$ is the event that the top pillar reaches height $h$, then [5, Thm. 3.8] (in the case of the Ising model, [12, Thm. 4.2]) proves that
@@
We now argue that Eq. 4.2, follows from a simplification of the same arguments of [5] that yielded Eqs. 4.3 and 4.4. The proof of Eq. 4.3 began by defining a cut point as a vertex $v\in\mathcal{P}^{\mathsf{top}}_{\mathsf{o}}$ with no other vertices in $\mathcal{P}^{\mathsf{top}}_{\mathsf{o}}$ having the same height, and such that removing the four vertical plaquettes surrounding $v$ cuts $\mathcal{C}$ into two 1-connected components, with the component above $v$ contained in $\mathcal{L}_{>\operatorname{ht}(v)}$. The portion of $\mathcal{C}$ above the lowest cut point was labeled the “spine” $\mathcal{S}_{\mathsf{o}}$. There was then a map which straightens the spine when it is too large, and deletes walls in the rest of $\mathcal{C}$ when they are too close to the spine. Doing so deletes the portion of $\mathcal{P}^{\mathsf{top}}_{\mathsf{o}}$ below the spine, so the map finishes by appending a column to rejoin the spine to the rest of $\mathcal{C}$. The proof follows the same way, treating all of $\mathcal{C}$ as the top pillar. Namely, one can define a cut point and the spine of the component $\mathcal{C}$, and apply the same map to the spine portion. Since under $\pi^{\mathsf{w}}_{n}$ there is no need to preserve the disconnection event ${\mathfrak{D}}_{n,n}$ (recall Eq. 2.4) as there was under $\bar{\pi}^{\mathsf{dob}}_{n}$, the remainder $\mathcal{C}\setminus\mathcal{S}_{\mathsf{o}}$ is handled by the simpler Peierls map which deletes it entirely and replaces it by a column of height $\operatorname{ht}(v)$ up from $\mathsf{o}$. Finally, Eq. 4.4 was shown by observing that when a portion of the spine is straightened, it becomes a column of vertices connected by open edges, and hence under the FK–Potts coupling they are all the same color. Clearly this observation holds regardless of whether the joint measure is ${\mathbf{P}}_{n}^{\mathsf{red}}$ or ${\mathbf{P}}_{n}^{\mathsf{dob}}$.
@@
(where we use the shorthand notation $a=b(1\pm\varepsilon_{\beta})$ to mean $a\in[b(1-\varepsilon_{\beta}),b(1+\varepsilon_{\beta})]$). We begin with (4.5). It is easy to see (by a Peierls argument) that $\pi^{\mathsf{w}}_{n}(\mathcal{E}^{\mathsf{w}})$ is dominated by (i.e., $1-\varepsilon_{\beta}$ of its mass is on) the event that $\mathcal{C}$ is just the six plaquettes surrounding $\mathsf{o}$, which happens with probability $qe^{-6\beta}(1\pm\varepsilon_{\beta})$. Moreover, by rigidity of $\mathcal{I}_{\mathsf{full}}$ (Theorem 2.5)
@@
We turn to establishing (4.6). First note that on $\mathcal{E}^{\mathsf{dob}}$, the event $\operatorname{ht}(\mathcal{P}_{\mathsf{o}}^{\mathsf{nonR}})\geq h$ is equivalent to $A_{\mathsf{o},h}$. Now, intending to utilize the FK–Potts coupling, we wish to first reveal in the random-cluster model the component $\mathcal{C}$. In order for the event $A_{\mathsf{o},h}$ to be possible, $\mathcal{C}$ must be a set of plaquettes that confines a chain of vertices connecting $\mathsf{o}$ to height $h$. Crucially, the events $\mathcal{E}^{\mathsf{w}}$ and $\mathcal{E}^{\mathsf{dob}}$ were defined so that the set of plaquettes $C^{*}$ such that $\pi^{\mathsf{w}}_{n}(\mathcal{C}^{*}=C^{*}\mid\mathcal{E}^{\mathsf{w}})$ and $\bar{\pi}^{\mathsf{dob}}_{n}(\mathcal{C}^{*}=C^{*}\mid\mathcal{E}^{\mathsf{dob}})$ both have positive probability is the same, let this set be denoted $\mathbf{C}^{*}$. In particular, $C^{*}\in\mathbf{C}^{*}$ if $C^{*}$ is a 1-connected component of plaquettes in $\mathbf{C}_{\vee}$ that forms part of a bounding surface in the sense that there can be no path in $\mathcal{L}_{>0}$ from $\mathsf{o}$ to $\partial\Lambda_{n,n}^{\prime}$ which does not cross a plaquette of $C^{*}\cup C_{1}$. Thus, we can express
@@
For the denominator, the event $\mathcal{E}^{\mathsf{dob}}$ is equivalent to asking that $\mathcal{C}^{*}\in\mathbf{C}^{*}$, and that the set of plaquettes $\mathcal{C}\setminus\mathcal{C}^{*}$ satisfies a set of criterion independent of $\mathcal{C}^{*}$. (Namely, that $\mathcal{C}\setminus\mathcal{C}^{*}$ contains $C_{1}$, is a subset of $\mathbf{C}_{\curlyvee}\cup C_{1}$, does not cut off $\mathsf{o}$ from $\partial_{\mathsf{blue}}\Lambda_{n,n}^{\prime}$, and combines with $\mathcal{C}^{*}$ such that $\mathcal{C}$ is an interface satisfying ${\mathfrak{D}}_{n,n}$. This last event appears as though the choice of $\mathcal{C}\setminus\mathcal{C}^{*}$ affects the possible choices of $\mathcal{C}^{*}$, but the cut-point at $\mathsf{o}$ means it does not.) Exposing $\mathcal{C}\setminus\mathcal{C}^{*}$ reveals a set of dual-to-open plaquettes around it by maximality. Let $F^{\mathsf{dob}}$ be the union of these plaquettes with $\mathcal{C}\setminus\mathcal{C}^{*}$. Noting that the separation event ${\mathfrak{D}}_{n,n}$ is included in $\mathcal{E}^{\mathsf{dob}}$, we can write $\bar{\pi}^{\mathsf{dob}}_{n}(\cdot\mid\mathcal{E}^{\mathsf{dob}})$ as $\mathbb{E}_{\mathcal{C}\setminus\mathcal{C}^{*}}[\pi^{\gamma}_{\Lambda_{n,n}^{\prime}\setminus F^{\mathsf{dob}}}(\cdot\mid\mathcal{C}^{*}\in\mathbf{C}^{*})\mid\mathcal{E}^{\mathsf{dob}}]$ where $\gamma$ is the boundary condition induced by revealing $\mathcal{C}\setminus\mathcal{C}^{*}$.